### v1

# Atomistic modeling of the channeling process with account for ionizing collisions: Preliminary results

Abstract
We present preliminary results of simulations of the channeling process with account for the ionizing collisions that accompany the passage of a charged projectile in a crystalline medium. The simulations have been performed by means of the MBN Explorer software package. taking into account that ionizing collisions are random, fast and local they can be incorporated into the classical molecular dynamics framework according to their probabilities. The description of this methodology is given in the paper. The case studies presented refer to channeling of 195, 270 & 855 MeV electrons and of 270 & 500 MeV positrons in single diamond crystals along the (111) (electrons) and (110) (positrons) planar directions.
1 Introduction
Interaction of high-energy charged particles with crystals is sensitive to the orientation of the incoming beam with respect to to the main crystallographic directions (planar and axial) of the target. Projectiles, incident at small angles to the crystal planes (or axes) experience specific channeling motion following the planar (axial) direction experiencing collective action of the electrostatic fields of the lattice atoms [1].
The study of the channeling process of ultra-relativistic particles in oriented crystals has emerged as a wide-ranging field of research [2, 3, 4, 5]. Various applications have been suggested including beam steering [6, 7, 8], collimation [9], focusing [10], and extraction [2].
Oriented crystals of different geometries (single, bent, periodically bent) exposed to the beams of ultra-relativistic electrons and positrons can potentially serve as novel intensive gamma-ray crystal-based light sources (CLS) operating in the MeV-GeV photon energy range [5, 11, 12]. Practical realization of CLSs is a subject of the running Horizon Europe EIC-Pathfinder-2021 project TECHNO-CLS [13]. This Project presents a high-risk/high-gain science-towards-technology breakthrough research aimed at further development of technologies for crystal samples preparation, extensive experimental program that includes design and manipulation of particle beams, detection and characterization of the radiation, as well as theoretical analysis and advanced computational modelling for the characterisation of CLS.
In recent years, advanced atomistic computational modeling of the channeling process and the process of the photon emission as well as of other related phenomena beyond the continuous potential framework has been carried out using the multi-purpose computer package MBN Explorer [14, 15] and the special multi-task software toolkit with graphical user interface MBN Studio [16]. The computation accounts for the interaction of projectiles with separate atoms of the environments, whereas many different interatomic potentials implemented in MBN Explorer support rigorous simulations of various media [17, 18]. For simulations of the channeling and related phenomena, an additional module has been incorporated into MBN Explorer to compute the motion of relativistic projectiles along with dynamical simulations of the environments. This methodology, called relativistic MD, has been extensively applied to simulate passage of ultra-relativistic charged particles in oriented crystals accompanied by emission of intensive radiation. A comprehensive description of the multiscale all-atom relativistic molecular dynamics approach as well as a number of case studies related to modeling channeling and photon emission by ultra-relativistic projectiles are presented in a review article [19].
Additional feature that has been implemented in MBN Explorer and is discussed in Section 2 below concerns the impact of the ionizing collisions on the channeling properties of ultra-relativistic charged particles. The events of inelastic scattering of a projectile particle from individual atoms lead to a random change in the direction of the particle’s velocity resulting in an additional mechanism of the increase in its transverse energy. Rigorous treatment of the ionization process can only be achieved by means of quantum mechanics. However, taking into account that such events are random, fast and local they can be incorporated into the classical molecular dynamics framework according to their probabilities. Similar methodology has been already implemented in MBN Explorer within the framework of the irradiation-driven molecular dynamics [26]. Preliminary results obtained for several related case studies are presented in Section 3.
2 Methodology
The MBN Explorer package [14, 17] implements the method of relativistic classical molecular dynamics [15] to model the motion an ultra-relativistic particle in an external field or/and in an atomic environment. The code integrates the relativistic equations of motion of a particle of mass $m$, charge $q$ and energy $\varepsilon$ written in the following form (see, e.g., [20]):
$\left\{\begin{array}[]{l}\dot{{\bf v}}=\displaystyle{{\bf F}-\bm{\beta}\left({\bf F}\cdot\bm{\beta}\right)\over m\gamma}\\ \dot{{\bf r}}=\textbf{v}\\ \end{array}\right.\ ,$ (3)
where ${\bf r}={\bf r}(t)$ and ${\bf v}={\bf v}(t)$ are instantaneous position vector and velocity of the particle, $\bm{\beta}={\bf v}/c$, $\gamma=\varepsilon/mc^{2}$ is the Lorentz factor, and $c$ is the speed of light.
In a general case, the force ${\bf F}$ acting on the projectile is the sum of two terms:
${\bf F}={\bf F}_{\rm em}+{\bf F}_{\rm rd}\,.$ (4)
Here ${\bf F}_{\rm em}$ denotes the Lorentz force due to (i) electrostatic field ${\bf E}$ created by atoms in the environment and/or by external charges, and (ii) external magnetic field ${\bf B}$:
${\bf F}_{\rm em}=q\left({\bf E}+\bm{\beta}\times{\bf B}\right).$ (5)
The second term in (4), ${\bf F}_{\rm rd}$, is a radiation damping force that accounts for the reaction of the emitted radiation on the motion of the charged particle, see, e.g. [20, 21]. Its action leads to a gradual decrease of the particle’s energy. For ultra-relativistic projectiles of very high energies (tens of GeV and above) this force becomes quite noticeable, so that it must be accounted for to properly simulate the motion, see recent experiments [22, 23, 24]. The methodology that has been developed to incorporate ${\bf F}_{\rm rd}$ into the relativistic molecular dynamics framework implemented in the MBN Explorer package is described in Ref.[25].
Applied to a charged particle’s passage through an atomic environment, the system (3) describes the classical motion of the particle in an electrostatic field of the medium’s atoms. In this case, both terms on the right-hand side of Eq. (5) are dependent on the field’s intensity which is is calculated as ${\bf E}=-\bm{\nabla}\phi({\bf r})$ with $\phi({{\bf r}})$ standing for the field’s potential equal to the sum of potentials of individual atoms
$\phi({\bf r})=\sum_{j}\phi_{\rm at}\left(\left|{\bf r}-{\bf r}_{j}\right|\right),$ (6)
where ${\bf r}_{j}$ denotes the position vector of the nucleus of the $j$th atom. For a neutral atom, the potential $\phi({\bf r})$ is a rapidly decreasing function at the distances larger than the average atomic radius. Therefore, at each step of integration of Eqs. (3), the sum in (6) can account only for the atoms inside the sphere of a specified cut-off radius $\rho$ with the center at ${\bf r}$. Typically, $\rho$ is chosen to be order of magnitude larger than the average atomic radius. The search for such atoms is carried out by means of the linked cell algorithm implemented in the MBN Explorer package. In the course of particle’s passage through a medium, the environment is generated dynamically so that the simulation box follows the moving particle. For a crystalline environment, the positions of the atoms are generated accounting for random displacement from the nodes due to thermal vibrations. More detailed description on the algorithms used to compute the trajectories at various scales, including macroscopically large ones, are presented in Refs. [14, 15, 17, 19, 25].
Classical equations of motion (3) do not account for random events of inelastic scattering of a projectile particle from individual atoms leading to the atom’s excitation or ionization. These events results in the two following changes in the particle motion:
  • Random decrease in the projectile energy (so-called, ionization losses).
  • Random change in the direction of the projectile motion characterized by two scattering angles $\theta$ and $\phi$ measured with respect to the instant velocity ${\bf v}$.
Rigorous treatment of the inelastic collision evens can only be achieved by means of quantum mechanics. However, taking into account that such events are random, fast and local they can be incorporated into the classical mechanics framework according to their probabilities. This approach has been already implemented in MBN Explorer within the framework of irradiation-driven molecular dynamics [26]. Similar methodology, described below, has been realized in the code to account for the ionizing collisions. Its implementation followed the scheme presented in Ref. [27].
For an ultra-relativistic particle passing through a medium, the probability ${\rm d}^{2}W$ of the relative energy loss ${\rm d}\Delta={\rm d}\varepsilon/\varepsilon$ over the path ${\rm d}s\approx c{\rm d}t$ due to ionizing collision with quasi-free electrons can be calculated using the following expression (see, e.g., Eqs. (1.7)-(1.7b) in [28] and §33.2.7 in [29]):
${{\rm d}^{2}W\over{\rm d}\Delta\,{\rm d}s}=2\pi r_{0}^{2}q^{2}\,{n_{{\rm e}}({\bf r})\over\gamma\Delta^{2}}\,.$ (7)
Here $r_{0}=2.818\times 10^{-13}$ stands for the classical electron radius and $n_{\mathrm{e}}(\textbf{r})$ is the local volume density of electrons. In the points ${\bf r}\neq{\bf r}_{j}$, i.e. not at the nuclei, the density is related to the potential (6) via the Poisson equation:
$n_{\mathrm{e}}(\textbf{r})={1\over 4\pi e}\bm{\nabla}^{2}\sum_{j}\phi_{\rm at}\left(\left|{\bf r}-{\bf r}_{j}\right|\right)\,.$ (8)
In a single ionizing collision, the value of relative energy loss $\Delta=(\varepsilon-\varepsilon^{\prime})/\varepsilon$, where $\varepsilon$ and $\varepsilon^{\prime}<\varepsilon$ are the particle’s energies before and after the collision, lies within the interval
$\Delta_{\min}\leq\Delta\leq\Delta_{\max},$ (9)
$\Delta_{\min}=I/\varepsilon,\qquad\Delta_{\max}\approx\left\{\begin{array}[]{ll}\displaystyle{2\gamma\xi\over 1+2\gamma\xi}&\mbox{for a heavy particle}\\ 1&\mbox{for a positron}\\ 0.5&\mbox{for an electron}\end{array}\right.$ (13)
where $\xi$ is the electron-to-projectile mass ratio and $I$ denotes the atomic ionization potential, which can be estimated as $I=16Z^{0.9}$ eV within the Thomas-Fermi model ($Z$ is the atomic number).
As written, the distribution (7) holds for $\Delta\ll 1$. However, since a probability of collisions with $\Delta\gtrsim 1$ is negligibly small, Eq. (7) can be applied to the whole range of the $\Delta$ values.
At each step $\Delta t$ of integration of Eqs. (3) an ionizing collision is treated as a probabilistic event. Once it occurs and the value $\Delta$ is determined, one calculates a round scattering angle $\theta$ measured with respect to the instant velocity ${\bf v}(t)$:
$\cos\theta=\sqrt{\gamma+1\over\gamma-1}\sqrt{1-\Delta-\gamma^{-1}\over 1-\Delta+\gamma^{-1}}-{\Delta(\xi-1)\over\sqrt{\gamma^{2}-1}\sqrt{(1-\Delta)^{2}-\gamma^{-2}}}$ (14)
The second scattering angle $\phi$ is not restricted by any kinematic relations, and its value is generated as a uniform random deviate between 00 and $2\pi$.
The algorithm of random generation of the $\Delta$ and $\theta$ values is as follows [27].
The two-fold probability (7) is normalized as follows
$\int_{0}^{L_{\rm ion}}\,\int_{\Delta_{\min}}^{\Delta_{\max}}\,{{\rm d}^{2}W\over{\rm d}\Delta\,{\rm d}{\rm d}s}\,{\rm d}\Delta\,{\rm d}s=1$ (15)
where $L_{\rm ion}$ is a spatial path over which the probability of an ionizing collision with an arbitrary energy transfer is equal to one:
$L_{\rm ion}^{-1}={2\pi r_{0}^{2}q^{2}\,n_{\mathrm{e}}(\textbf{r})\over\gamma}\,{1\over\Delta_{0}},\qquad\Delta_{0}={\Delta_{\max}\,\Delta_{\min}\over\Delta_{\max}-\Delta_{\min}}\approx\Delta_{\min}\,.$ (16)
Therefore, the probability of an ionizing collision with the energy transfer $\Delta$ to occur over the path $\Delta s=c\Delta t$ is given by
${\rm d}W={\Delta s\over L_{\rm ion}}\,P(\Delta)\,{\rm d}\Delta\,,\quad P(\Delta)={\Delta_{0}\over\Delta^{2}}$ (17)
where the factor $P(\Delta)\,{\rm d}\Delta$ stands the normalized probability of the relative energy loss between $\Delta$ and $\Delta+{\rm d}\Delta$.
To simulate the probability of an ionizing collision to take place, a uniform random deviate $x\in[0,1]$ is generated and compared with $\Delta s/L_{\rm ion}$. If $x\leq\Delta s/L_{\rm ion}$ then the collision happens and another random deviate $\Delta$ is generated and used in Eq. (14) to calculate $\theta$.
The values of $\Delta$, $\theta$ and $\phi$ obtained are used further to modify the velocity and energy of the projectile at the start of the next integration step.
To account for or to disregard the ionizing collisions the following schemes are implemented in MBN Explorer:
  • the local electron density $n_{\mathrm{e}}(\textbf{r})$ is calculated using Eq. (8) where atomic potentials $\phi_{\rm at}$ are considered within the frameworks of the Molière [30] or Pacios [31] approximations.
  • the local density is substituted with the average electron density in the amorphous medium: $n_{\mathrm{e}}(\textbf{r})=Zn_{\rm at}$ where $n_{\rm at}$ stands for the volume density of atoms in the amorphous environment,
  • to disregard the ionizing collisions one sets $n_{\mathrm{e}}(\textbf{r})\equiv 0$.
3 Case studies: Preliminary results
The case studies presented below refer to the channeling of $\varepsilon=195,270,855$ MeV electrons and $\varepsilon=270,500$ MeV positrons in 200 and 400 microns thick single diamond crystals. For electrons, the aforementioned energy range corresponds that available at the Mainz Mikrotron (MAMI) facility. A number of channeling experiments with the electron beam have been carried out at MAMI with various crystal targets during the last one and a half decade [32, 33, 6, 34, 35, 36]. More recently, within the framework of the TECHNO-CLS project [13], a new 500 MeV positron beam line has been designed at MAMI [37] and the first experimental tests with the positron beam have been carried out [38, 39].
The aim of the present study was to quantify the impact of the ionizing collisions on the channeling efficiency of projectiles of different charge and energy passing through comparatively thick crystalline targets. For doing this we have followed general methodology implemented in MBN Explorer to generate particles’ trajectories in a crystalline environment that accounts for randomness in sampling the incoming projectiles (the key factors here are beam size, divergence and direction with respect to the crystal orientation) as well as in displacement of the lattice atoms from the nodal positions due to thermal vibrations (for more details see Refs. [15, 19]). As a result, each trajectory corresponds to a unique crystalline environment and, therefore, all simulated trajectories are statistically independent and can be analyzed further to quantify the channeling process.
For each case study approximately $N_{0}=1.2\times 10^{4}$ of trajectories have been simulated and analyzed. The initial analysis has been carried out in terms of the two channeling fractions: $f_{\rm ch,0}(z)=N_{\rm ch,0}(z)/N_{0}$ and $f_{\rm ch}(z)=N_{\rm ch}(z)/N_{0}$ [19]. Here $N_{\rm ch,0}(z)$ denotes for the number of particles that were accepted to the channeling mode at the crystal entrance ($z=0$) and move in the channel up to distance $z$ where they dechannel. The ratio ${\cal A}=N_{\rm ch,0}(0)/N_{0}$ defines the fraction of the incoming particles accepted at the entrance. The quantity $N_{\rm ch}(z)$ stands for the total number of particles which move in the channeling mode at the distance $z$.
At comparatively large distances $z$ from the entrance point the fractions $f_{\rm ch,0}(z)$ decreases following the exponential decay law $\exp(-z/L_{\rm d})$, where $L_{\rm d}$ is the dechanneling length [1, 40]. The fraction $f_{\rm ch}(z)$ can be approximated by means of the fitting formula [19]:
$f_{\rm ch}(z)\approx{\rm erf}\left(\beta\sqrt{\varepsilon\over z}\right)$ (18)
where ${\rm erf}(.)$ stands for the error function.
The parameters $L_{\rm d}$ and $\beta$ can be defined by comparing the fitting formulae with the dependencies obtained in the simulations.
Figure 1: Channeling fractions $f_{{\rm ch},0}$ (dashed lines) and $f_{\rm ch}$ (solid lines) versus penetration distance $z$ for $\varepsilon=195,270$ and 855 MeV electrons in 200 $\mu$m thick diamond(111) crystal. Curves without symbols show the results obtained without accounting for the ionizing collisions; curves with symbols correspond to the dependencies calculated using the Molière potential to simulate the probability (7).
[…]
The main conclusion, which one can draw comparing the dependencies presented in Fig. 1, is that for electrons the ionizing collisions can be disregarded in the quantitative description of the channeling motion. This can be understood recalling that negatively charged particles experience channeling motion in the vicinity of the atomic planes. Therefore, they experience rather close collisions (i.e. with comparatively small impact parameter) with the crystal atoms. At these impact parameters the contribution of the inelastic channels to the total cross section is much smaller than of the elastic collision, see, for example, [41].
Figure 2: Same as in Fig. 2 but for 270 and 500 MeV positrons in 400 $\mu$m thick diamond(110) crystal.
The dependencies $f_{{\rm ch},0}$ (dashed lines) and $f_{\rm ch}$ (solid lines) obtained in the simulations of the 270 and 500 MeV positron channeling in a 400 $\mu$m thick single diamond(110) crystal performed without (no symbols) and with (open circles and squares) account for the ionizing collisions are shown in Fig. 2. In this case, at large penetration distances the values of both fractions calculated with account for the ionizing collisions are about 15 % (for $\varepsilon=270$ MeV) and 10 % (for 500 MeV) lower than those obtained ignoring the ionization. It is not surprising though that the impact of the ionizing collisions on the channeling efficiency of positrons is much stronger that for electrons. Positively charged particles effectively channel in the spatial regions between atomic planes, therefore, statistically they experience much more distant collisions with atoms than electrons. At large impact parameters the inelastic channels become dominant in the total cross section of the collision between an ultra-relativistic projectile and an atom. As a result, the increase in the scattering angle (and, consequently, in the transverse energy) occurs at higher rate leading to faster decrease in the channeling fractions with $z$.
4 Conclusions
We have reported the results of preliminary statistical analysis of the role which inelastic collisions with the crystal atoms play in the process of planar channeling of ultra-relativistic electrons and positrons of sub-GeV energies. The simulations have been performed by means of the MBN Explorer software package within the framework of relativistic classical molecular dynamics. The events of inelastic scattering of a projectile particle from crystal atoms have been incorporated into the classical framework according to their probabilities. On the basis of the simulations performed the dependence of the channeling fractions on the penetration distance has been obtained for several eelectron and positron beam energies.
More consistent analysis of the impact of the ionizing collisions on the dynamics of channeling as well as on the characteristics of the radiation emitted by channeling particles will be published elsehwere.
Acknowledgements
We acknowledge support by the European Commission through the N-LIGHT Project within the H2020-MSCA-RISE-2019 call (GA 872196) and the EIC Pathfinder Project TECHNO-CLS (Project No. 101046458). We also acknowledge the Frankfurt Center for Scientific Computing (CSC) for providing computer facilities.